Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night Report Page - 20121107

The previous night we noticed a possible misalignment of grism13 (~13pixel of displacement along the CCD in the dispersion direction), so we sent a new setup form in order to ask a check of the wheels elements. For this, we had to get a new set of calibration (arcs and dome flat fields) in addition to those obtained at the beginning of the subrun.

During the afternoon, we fixed the problem with the spectropolarimetric OBs and took new calibrations.


Cirrus and thin clouds at the beginning of the night. Clear sky at the end. No major problem apart for the spectropolarimetry.

Seeing 0.8-1"


EFOSC2
  Skyflats g
  UT23:45  HD13588 unpol STD
  2D spectra looked too faint, possibly because of the cirrus. 
  We increased the exposure time of the OB up to 12 times than what foreseen, keeping getting a very low signal. 
  TOs suggested an even more long exposure time for the gm13 setup.
  Since the planned 3h of exposures seemed to be not enough to get useful spectropolarimetric data of SN2012fr this night,
  we decided to skip it for this subrun. CHECK carefully exposure times for the next runs.
  THE DAY AFTER: the problem was most probably in the binning of the through-slit image acquired in the afternoon, that gave
  us the wrong position on the slit for the source 


  UT00:45  Feige110-STD  g11,g13,g16
  UT01:02  SN2012dy  g16
  UT02:20  SN2009ip  g11,g16, UBVRi
  UT03:20  LSQ12dyw  UVBRi,g13
  UT04:52  OGLE-2012-SN-027 CLASS [SNID: SNIa z=0.24 at max ??]
  UT05:28  LSQ12gak  CLASS [Likely too faint: No SN identified in the acquisition image]
  UT05:46  SN2012fr  UBVRi,g11,g16
  UT06:07  CHASE_ESO-9015   CLASS [SNID: SNIIP z=0.009 136d]
  UT06:22  OGLE-2012-SN-029 CLASS [no SN identified in the acquistion image, too faint]
  UT06:30  EG21-STD   g11,g13,g16

SOFI (UT06:40)---
  UT06:50 SN2012fr  JHK and BG,RG
  UT08:07 Hip012167-Telluric BG,RG
  UT08:13 LTT3218-Flux BG,RG
Comments